OpenStudy (anonymous):

PLEASE HELP!!!!!!!!!!!!!!!!!!!! Three years ago, Andy invested $7,000 in an account that earns 10% interest compounded annually. The equation y = 7,000(1.10)t describes the balance in the account, where t is time in years. Andy made no additional deposits and no withdrawals. How much is in Andy’s account at this time? Round to the nearest dollar.

OpenStudy (mertsj):

Replace t with 3 and evaluate.

OpenStudy (anonymous):

23100???

OpenStudy (mertsj):

Isn't that t an exponent?

OpenStudy (mertsj):

Your answer is way too big.

OpenStudy (anonymous):

THE T behind (1.10) is

OpenStudy (mertsj):

So here is your equation: \[y=7000(1.10)^3\]

OpenStudy (mertsj):

What is 1.10 raised to the third power?

OpenStudy (anonymous):

where did you get the 3 from

OpenStudy (mertsj):

Three years ago Andy invested.... How much is in Andy’s account at this time?

OpenStudy (mertsj):

What do you think t is?

OpenStudy (anonymous):

nvm

OpenStudy (anonymous):

my notes had something different

OpenStudy (mertsj):

Like what?

OpenStudy (anonymous):

Like it tells you the words if they are in the word problem you have to use it like annually= one per year, bi-monthly=6 times per year You would use the number in your equation

OpenStudy (mertsj):

The formula is: \[A=A _{0}(1+\frac{r}{n})^{nt}\]

OpenStudy (anonymous):

loook

OpenStudy (mertsj):

A is the final amount after t years A_0 is the beginning amount. r is the interest rate n is the number of compounding periods in 1 year

OpenStudy (anonymous):

did u see the website thats my notes

OpenStudy (mertsj):

Yes. But I can only see the first page because I cannot log in.

OpenStudy (mertsj):

Did you see the equation I posted?

OpenStudy (anonymous):

yes i did . You have to use a number 1 in the equation because it says the word annually right ?

OpenStudy (mertsj):

In your problem the beginning amount is 7000 n = 1 compounding period each year r = .10 t = 3 So your problem becomes: \[A=7000(1+\frac{.10}{1})^{1(3)}\]

OpenStudy (mertsj):

Which is the same as: \[A=7000(1.10)^3\]

OpenStudy (mertsj):

Which is: \[A=7000(1.331)\]

OpenStudy (mertsj):

Which is: \[A=$9317\]

OpenStudy (anonymous):

theres the same example ^^ with the word annually

OpenStudy (anonymous):

did u see it ??

OpenStudy (mertsj):

Yep. Exactly what I just posted.

OpenStudy (mertsj):

Did you see what I posted?

OpenStudy (anonymous):

Im going over it & i understand it thanks! I hope its right

OpenStudy (mertsj):

It is right.

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

The value of Mr. Cox’s car x years after its purchase is given by the function, V(x) = 20,000(0.85)x. Approximately, what was the value of Mr. Cox’s car 5 years after its purchase? Round to the nearest dollar.

OpenStudy (anonymous):

How would u do this ?

OpenStudy (mertsj):

I would make x the exponent and replace it with 5

OpenStudy (anonymous):

(0.85)^x

OpenStudy (anonymous):

so then you would just multiply it ?

OpenStudy (mertsj):

\[V(5)=20000(.85)^5=20000(.4437)=8874.10\]

OpenStudy (mertsj):

$8874 to the nearest dollar.

OpenStudy (anonymous):

thats the answer ??

OpenStudy (anonymous):

if it is it it rounded

OpenStudy (mertsj):

yep

OpenStudy (anonymous):

thanks

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

can u help me on this one too ? A city’s current population is 1,000,000 people. It is growing at a rate of 3.5% per year. The equation P = 1,000,000(1.035)x models the city’s population growth where x is the number of years from the current year. In approximately how many years will the population be 1,200,000? Round to the nearest tenth.

OpenStudy (mertsj):

Replace P with 1,200,000 and solve for x.

OpenStudy (anonymous):

1,200,000=1,000,000(1.035)x ?

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

would you multiply 1,000,000(1.035)x

OpenStudy (anonymous):

opps its supposed to be (1.035)^x

OpenStudy (mertsj):

I know.

OpenStudy (anonymous):

.8625 ??

OpenStudy (anonymous):

is this right

OpenStudy (anonymous):

you there ???

OpenStudy (mertsj):

I haven't done it yet.

OpenStudy (mertsj):

\[1.2=1.035^x\] \[\log_{10}1.2=x \log_{10}1.035 \] \[\frac{\log_{10} 1.2}{\log_{10}1.035}=x \] x=5.3 years

OpenStudy (anonymous):

do i round ?

OpenStudy (mertsj):

What does the problem say to do?

OpenStudy (anonymous):

round to the nearest tenth so its 10 !!

OpenStudy (anonymous):

I tried 1 and it was wrong .

OpenStudy (anonymous):

or is it 5 ?

OpenStudy (mertsj):

The nearest TENTH not the nearest 10. Do you know place value?

OpenStudy (anonymous):

yes